# InterCLIP-MEP: Interactive CLIP and Memory-Enhanced Predictor for Multi-modal Sarcasm Detection

### v3

@@ Introduction
Sarcasm, with its inherent subtlety and complexity, plays a significant role in human communication by often conveying irony, mockery, or hidden intentions (Muecke 1982; Gibbs and O’Brien 1991; Gibbs and Colston 2007).
Figure 1: Comparative analysis of multi-modal sarcasm detection pipelines. Panels (a), (b), and (c) depict the pipeline comparisons between our InterCLIP-MEP method and other existing methods. Panel (d) utilizes GradCAM to visually demonstrate the superior ability of the InterCLIP backbone over the traditional CLIP backbone in identifying multi-modal sarcasm cues, which are marked with red boxes.
The automatic detection of sarcasm from text has been a significant research focus, aiding tasks such as sentiment analysis and intent mining (Pang, Lee et al. 2008; Tsur, Davidov, and Rappoport 2010; Bouazizi and Ohtsuki 2015). With the rapid development of social media platforms like Twitter and Reddit, users frequently use text-image combinations to convey their messages. Therefore, the need for multi-modal sarcasm detection has become increasingly apparent, which presents significant challenges, requiring the system to capture and comprehend the intricate interplay between textual and visual information to identify sarcasm cues.
@@
As shown in Figure 1(a) and Figure 1(b), some existing methods (Xu, Zeng, and Mao 2020; Pan et al. 2020; Liang et al. 2021, 2022; Wen, Jia, and Yang 2023; Tian et al. 2023; Wei et al. 2024) use separate uni-modal encoders, such as BERT (Devlin et al. 2019) for text and ViT (Dosovitskiy et al. 2020) for image, as the backbone to encode samples. This approach makes it difficult for the model to capture the subtle interactive information between text and image. Although Qin et al. (2023) utilized a multi-modal pre-trained backbone, CLIP (Radford et al. 2021), to encode samples and demonstrated notable results on the MMSD2.0 benchmark, the original CLIP only aligns text and images directly, lacking the nuanced semantic space required for sarcasm detection. Furthermore, as depicted in Figure 1(a), some approaches (Xu, Zeng, and Mao 2020; Pan et al. 2020; Liang et al. 2021, 2022; Wen, Jia, and Yang 2023; Tian et al. 2023) directly classify sarcastic samples based on their representations. This often leads to high prediction entropy and erroneous outcomes, especially for hard cases with greater uncertainty. To the best of our knowledge, only Wei et al. (2024) considers leveraging historical knowledge of samples to aid in detection. As shown in Figure 1(b), they constructed a static semantic space that stores features of all training samples. When identifying sarcastic samples, they query $k$-nearest neighbors in this semantic space and use a voting mechanism to derive the final predictions. However, relying heavily on a static semantic space may limit its ability to adapt to various test samples.
@@
In this paper, building on these insights, we propose InterCLIP-MEP, a novel framework designed to enhance the reliability of multi-modal sarcasm detection. As shown in Figure 1(c), our InterCLIP-MEP introduces a refined variant of CLIP, Interactive CLIP (InterCLIP), as the backbone to enhance sample encoding for detecting multi-modal sarcasm. InterCLIP enhances the capture of interactive information between text and image by embedding representations from one modality into the encoder of the other, leading to a more nuanced understanding of multi-modal sarcasm cues. In Figure 1(d), we utilized GradCAM (Selvaraju et al. 2016) visualization to compare the effectiveness of our InterCLIP backbone against the CLIP backbone in detecting multi-modal sarcasm cues. During inference, InterCLIP-MEP introduces a Memory-Enhanced Predictor (MEP) to build dynamically a non-parametric classifier from valuable historical features to identify multi-modal sarcasm. Specifically, we design an efficient training strategy to adapt InterCLIP for MEP. Based on the sample representations encoded by InterCLIP, we train a classification module to assign pseudo-labels to samples. Simultaneously, we train a projection module to map samples into a latent space. In this latent space, same-class samples are pulled closer together, while different-class samples are pushed further apart. Additionally, we fine-tune the weights of the self-attention modules in InterCLIP’s text and vision encoders using the low-rank adaptation (LoRA) technique (Hu et al. 2022), building upon the original CLIP weights. This approach allows for more efficient adaptation without significantly increasing the computational burden. Building on InterCLIP and the training strategy, MEP leverages historical knowledge of valuable test samples to identify multi-modal sarcasm. It maintains a dynamic, fixed-length dual-channel memory, with one channel storing historical information of non-sarcastic samples and the other storing information of sarcastic samples. MEP then utilizes this memory as a non-parametric classifier to compute the final prediction for each sample, offering more robust and reliable multi-modal sarcasm detection.
@@
Figure 2: Overview of our framework. The left part of the figure shows the overall structure and training process of Interactive-CLIP. The right part illustrates the process of MEP identifying sarcastic samples.
@@
Multi-modal sarcasm detection. Early work in sarcasm detection primarily focused on textual data (Bouazizi and Ohtsuki 2015; Amir et al. 2016; Baziotis et al. 2018). With the rise of social media, detecting sarcasm from text-image data has become increasingly challenging, leading to the development of several multi-modal approaches. Schifanella et al. (2016) were among the first to use multi-modal posts from social platforms to identify sarcasm clues from both text and images. Cai, Cai, and Wan (2019) introduced the MMSD benchmark for multi-modal sarcasm detection, demonstrating its effectiveness with a hierarchical fusion model incorporating image attributes. This benchmark has been widely used in subsequent studies (Xu, Zeng, and Mao 2020; Pan et al. 2020; Liang et al. 2021, 2022; Liu, Wang, and Li 2022; Qin et al. 2023; Wen, Jia, and Yang 2023; Tian et al. 2023; Wei et al. 2024). However, MMSD has been proven to contain spurious cues that could bias models (Qin et al. 2023). To address this, Qin et al. (2023) introduced the MMSD2.0 benchmark, which removed these cues and corrected mislabeled samples. They observed a significant performance drop when re-evaluating state-of-the-art methods on MMSD2.0 and proposed the need for more stable multi-modal sarcasm detection methods. In this work, we introduce the InterCLIP-MEP framework towards more reliable multi-modal sarcasm detection.
@@
Mermory-enhanced prediction. Inspired by cognitive science (Stokes 2015; Baddeley 2000), memory has been introduced to enhance neural networks (Weston, Chopra, and Bordes 2014; Sukhbaatar et al. 2015). Several studies (Wu et al. 2018; Wen, Jia, and Yang 2023) have used memory mechanisms to improve model training, and some (Zhang et al. 2024; Wei et al. 2024) leverage memory to store historical knowledge, enhancing prediction accuracy. In this work, we introduce a memory-enhanced predictor for multi-modal sarcasm detection. In contrast to other methods, our memory dynamically updates during testing, utilizing relevant historical information for improved accuracy and robustness.
@@
$\mathbf{F}_{t}=\mathcal{T}(T)=\{h^{t}_{\textrm{bos}}(t_{\textrm{bos}}),\\ h^{t}_{1}(t_{1}),\dots,h^{t}_{n}(t_{n}),h^{t}_{\textrm{eos}}(t_{\textrm{eos}})\},$ (1)
@@
In particular, we condition only the top-$n$ self-attention layers of the text or vision encoder, where $n$ is a hyperparameter that will be analyzed in the experiment section. The structure of the conditioned self-attention layers is illustrated in Figure 3. We denote the input representations to the self-attention layers as $\mathbf{H}_{t/v}$, which are the output representations from the previous layer. The previous layer can either be conditioned or non-conditioned. Due to the dimensional mismatch between the embedded representations $\mathbf{F}_{v/t}$ and the corresponding encoder representation space, we introduce an adapting projection layer $\mathcal{F}_{t/v}$ to project $\mathbf{F}_{v/t}$ into the appropriate representation space.
@@
$\mathbf{F}_{v}^{{}^{\prime}}=\mathcal{F}_{t}(\mathbf{F}_{v}),\mathbf{F}_{v}\in\mathbb{R}^{{m}\times{d_{v}}},\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{{m}\times{d_{t}}},$ (4)
$\mathbf{H}^{{}^{\prime}}_{t}=\textrm{Attn}_{t}(\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}})_{\textrm{[:n]}},\\ \mathbf{H}_{t},\mathbf{H}^{{}^{\prime}}_{t}\in\mathbb{R}^{{n}\times{d_{t}}},\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{{(n+m)}\times{d_{t}}},$ (5)
$\mathbf{H}^{{}^{\prime\prime}}_{t}=\mathcal{H}_{t}(\mathbf{H}^{{}^{\prime}}_{t})+\mathcal{G}_{t}(\mathbf{H}^{{}^{\prime}}_{t})\cdot\textrm{tanh}(\beta_{t}),\mathbf{H}^{{}^{\prime\prime}}_{t}\in\mathbb{R}^{{n}\times{d_{t}}}.$ (6)
@@
Image ResNet* (He et al. 2015) 65.50 57.58 61.17 54.39
ViT* (Dosovitskiy et al. 2020) 72.02 69.72 65.26 74.83
@@
We compare the effectiveness of the InterCLIP-MEP framework against several uni-modal and multi-modal methods. For text modality methods, we compare with TextCNN (Kim 2014), Bi-LSTM (Graves and Schmidhuber 2005), SMSD (Xiong et al. 2019), and RoBERTa (Liu et al. 2019). For image modality methods, we compare with ResNet (He et al. 2015) and ViT (Dosovitskiy et al. 2020). For state-of-the-art multi-modal methods, we compare with the following:
@@
Figure 4: Hyperparameter study curves for InterCLIP-MEP w/ T2V. Figure 4(d) compares results with those from using only the classification module $\mathcal{F}_{c}$ for prediction.

### v4

@@ Introduction
Sarcasm, with its inherent subtlety and complexity, plays a significant role in human communication by often conveying irony, mockery, or hidden intentions (Muecke 1982; Gibbs and O’Brien 1991; Gibbs and Colston 2007). The automatic detection of sarcasm from text has been a significant research focus, aiding tasks such as sentiment analysis and intent mining (Pang, Lee et al. 2008; Tsur, Davidov, and Rappoport 2010; Bouazizi and Ohtsuki 2015). With the rapid development of social media platforms like Twitter and Reddit, users frequently use text-image combinations to convey their messages. Therefore, the need for multi-modal sarcasm detection has become increasingly apparent, which presents significant challenges, requiring the system to capture and comprehend the intricate interplay between textual and visual information to identify sarcasm cues.
Figure 1: An overview of the shortcomings of existing multi-modal sarcasm detection pipelines. In panels (a) and (b), we present two multi-modal sarcasm detection pipelines, with shortcomings indicated by a red question mark. In panel (c), we visually show an example of multi-modal sarcasm cues correctly or incorrectly recognized in a sarcasm sample.
@@
As shown in Figure 1(a) and Figure 1(b), some existing methods (Xu, Zeng, and Mao 2020; Pan et al. 2020; Liang et al. 2021, 2022; Wen, Jia, and Yang 2023; Tian et al. 2023; Wei et al. 2024) use dual unimodal encoders, such as BERT (Devlin et al. 2019) for text and ViT (Dosovitskiy et al. 2021) for image, as the backbone to encode samples. This approach makes it difficult for the model to capture the subtle interactive information between text and image. Although Qin et al. (2023) utilized a multi-modal pre-trained backbone, CLIP (Radford et al. 2021), to encode samples and demonstrated notable results on the MMSD2.0 benchmark, the original CLIP only aligns text and images directly, lacking the nuanced semantic space required for sarcasm detection. To illustrate the identification of multi-modal sarcasm cues, we present in Figure 1(c) an example of a sarcasm sample being correctly and incorrectly recognized. Furthermore, as depicted in Figure 1(a), some approaches (Xu, Zeng, and Mao 2020; Pan et al. 2020; Liang et al. 2021, 2022; Wen, Jia, and Yang 2023; Tian et al. 2023) directly classify sarcastic samples based on their representations. This often leads to high prediction entropy and erroneous outcomes, especially for hard cases with greater uncertainty. To the best of our knowledge, only Wei et al. (2024) considers leveraging historical knowledge of samples to aid in detection. As shown in Figure 1(b), they constructed a static semantic space that stores features of all training samples. When identifying sarcastic samples, they query $k$-nearest neighbors in this semantic space and use a voting mechanism to derive the final predictions. However, relying heavily on a static semantic space may limit its ability to adapt to various test samples.
@@
In this paper, building on these insights, we propose InterCLIP-MEP, a novel framework designed to enhance the reliability of multi-modal sarcasm detection. Our InterCLIP-MEP introduces a refined variant of CLIP, Interactive CLIP (InterCLIP), as the backbone to enhance sample encoding for detecting multi-modal sarcasm. InterCLIP enhances the capture of interactive information between text and image by embedding representations from one modality into the encoder of the other, leading to a more nuanced understanding of multi-modal sarcasm cues. During inference, Our InterCLIP-MEP introduces a Memory-Enhanced Predictor (MEP) to build dynamically a non-parametric classifier from valuable historical features to identify multi-modal sarcasm. Specifically, we design an efficient training strategy to adapt InterCLIP for MEP. Based on the sample representations encoded by InterCLIP, we train a classification module to assign pseudo-labels to samples. Simultaneously, we train a projection module to map samples into a latent space. In this latent space, same-class samples are pulled closer together, while different-class samples are pushed further apart. Additionally, we fine-tune the weights of the self-attention modules in InterCLIP’s text and vision encoders using the low-rank adaptation (LoRA) technique (Hu et al. 2022), building upon the original CLIP weights. This approach allows for more efficient adaptation without significantly increasing the computational burden. Building on InterCLIP and the training strategy, MEP leverages historical knowledge of valuable test samples to identify multi-modal sarcasm. It maintains a dynamic, fixed-length dual-channel memory, with one channel storing historical information of non-sarcastic samples and the other storing information of sarcastic samples. MEP then utilizes this memory as a non-parametric classifier to compute the final prediction for each sample, offering more robust and reliable multi-modal sarcasm detection.
@@
Figure 2: Overview of our framework. Training Interactive CLIP (InterCLIP). Vision and text representations are extracted using separate encoders and embedded into the top-n layers of the opposite modality’s encoder for interaction. The top-n layers are fine-tuned with LoRA, while the rest of the encoder remains frozen. Final vision and text representations are concatenated and used to train a classification module for sarcasm detection. A projection module is also trained to project representations into a latent space. Memory-Enhanced Predictor (MEP). During inference, InterCLIP generates interactive representations. The classification module assigns pseudo-labels, and the projection module provides projection features. MEP updates dynamic memory with these features and pseudo-labels. The final prediction is made by comparing the current sample’s projected feature with those in memory.
@@
Multi-modal sarcasm detection. Early work in sarcasm detection primarily focused on textual data (Bouazizi and Ohtsuki 2015; Amir et al. 2016; Baziotis et al. 2018). With the rise of social media, detecting sarcasm from text-image data has become increasingly challenging, leading to the development of several multi-modal approaches. Schifanella et al. (2016) were among the first to use multi-modal posts from social platforms to identify sarcasm cues from both text and images. Cai, Cai, and Wan (2019) introduced the MMSD benchmark for multi-modal sarcasm detection, demonstrating its effectiveness with a hierarchical fusion model incorporating image attributes. This benchmark has been widely used in subsequent studies (Xu, Zeng, and Mao 2020; Pan et al. 2020; Liang et al. 2021, 2022; Liu, Wang, and Li 2022; Qin et al. 2023; Wen, Jia, and Yang 2023; Tian et al. 2023; Wei et al. 2024). However, MMSD has been proven to contain spurious cues that could bias models (Qin et al. 2023). To address this, Qin et al. (2023) introduced the MMSD2.0 benchmark, which removed these cues and corrected mislabeled samples. They observed a significant performance drop when re-evaluating state-of-the-art methods on MMSD2.0 and proposed the need for more stable multi-modal sarcasm detection methods. In this work, we introduce the InterCLIP-MEP framework towards more reliable multi-modal sarcasm detection.
@@
Memory-enhanced prediction. Inspired by cognitive science (Stokes 2015; Baddeley 2000), memory has been introduced to enhance neural networks (Weston, Chopra, and Bordes 2014; Sukhbaatar et al. 2015). Several studies (Wu et al. 2018; Wen, Jia, and Yang 2023) have used memory mechanisms to improve model training, and some (Zhang et al. 2024; Wei et al. 2024) leverage memory to store historical knowledge, enhancing prediction accuracy. In this work, we introduce a memory-enhanced predictor for multi-modal sarcasm detection. In contrast to other methods, our memory dynamically updates during testing, utilizing relevant historical information for improved accuracy and robustness.
@@
$\mathbf{F}_{t}=\mathcal{T}(T)=\{h^{t}_{\textrm{bos}}(t_{\textrm{bos}}),h^{t}_{1}(t_{1}),\dots,h^{t}_{n}(t_{n}),h^{t}_{\textrm{eos}}(t_{\textrm{eos}})\},$ (1)
@@
To be specific, we condition only the top-$n$ self-attention layers of the text or vision encoder, where $n$ is a hyperparameter that will be analyzed in the experiment section. Figure 3 illustrates the structure of the conditioned self-attention layers. Given that the text and vision encoder in CLIP share a similar architecture, for simplicity, we denote the input representations to the self-attention layers of the text or vision encoder as $\mathbf{H}_{t/v}$, which are derived from the outputs of the previous layer. The previous layer can either be conditioned or non-conditioned. Due to the dimensional mismatch between the embedded representations $\mathbf{F}_{v/t}$ and the corresponding encoder representation space, we introduce an adapting projection layer $\mathcal{F}_{t/v}$ to project $\mathbf{F}_{v/t}$ into the appropriate representation space.
@@
$\begin{split}\mathbf{F}_{v}^{{}^{\prime}}&=\mathcal{F}_{t}(\mathbf{F}_{v}),\quad\mathbf{F}_{v}\in\mathbb{R}^{m\times d_{v}},\quad\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{m\times d_{t}},\\ \mathbf{H}^{{}^{\prime}}_{t}&=\textrm{Attn}_{t}(\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}})_{[:n]},\\ &\phantom{=}\mathbf{H}_{t},\mathbf{H}^{{}^{\prime}}_{t}\in\mathbb{R}^{n\times d_{t}},\quad\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{(n+m)\times d_{t}},\\ \mathbf{H}^{{}^{\prime\prime}}_{t}&=\mathcal{H}_{t}(\mathbf{H}^{{}^{\prime}}_{t})+\mathcal{G}_{t}(\mathbf{H}^{{}^{\prime}}_{t})\cdot\textrm{tanh}(\beta_{t}),\quad\mathbf{H}^{{}^{\prime\prime}}_{t}\in\mathbb{R}^{n\times d_{t}}.\end{split}$ (4)
@@
Image ResNet* (He et al. 2016) 65.50 57.58 61.17 54.39
ViT* (Dosovitskiy et al. 2021) 72.02 69.72 65.26 74.83
@@
We compare the effectiveness of the InterCLIP-MEP framework against several unimodal and multi-modal methods. For text modality methods, we compare with TextCNN (Kim 2014), Bi-LSTM (Graves and Schmidhuber 2005), SMSD (Xiong et al. 2019), and RoBERTa (Liu et al. 2019). For image modality methods, we compare with ResNet (He et al. 2016) and ViT (Dosovitskiy et al. 2021). For state-of-the-art multi-modal methods, we compare with the following:
@@
Figure 4: Hyperparameter study curves for InterCLIP-MEP w/ T2V. Panel (d) compares results with those from using only the classification module $\mathcal{F}_{c}$ for prediction.